OpenStudy (anonymous):

The surface area of this pyramid is one hundred eighty-nine square centimeters. The area of the square base is forty-nine square centimeters. What is the area of one triangular face of the lateral surface? Show the steps of your solution and explain your work. a square pyramid

OpenStudy (lxelle):

A square pyramid is formed by a square base and four triangular faces surface area of pyramid = area of square + 4(area of triangle) 189 = 49 + 4(area of triangle) area of triangle = (189 - 49) /4 = 140/4 = ???

OpenStudy (anonymous):

Thank you! :)

OpenStudy (lxelle):

No problem. :)

OpenStudy (anonymous):

@LXelle where did you get 140 Can you explain the steps to me?

OpenStudy (lxelle):

189-49 i got 140

OpenStudy (anonymous):

Do you think you may have made some (TYPO's) in your work above? Because it is not making sense... @LXelle

OpenStudy (lxelle):

surface area of pyramid = area of square + 4(area of triangle) 189 = 49 + 4(area of triangle) area of triangle = (189 - 49) /4 = 140/4 = ??? makes more sense now?

OpenStudy (anonymous):

Kind of... Thanks, I will try to put the pieces together...

OpenStudy (lxelle):

Which part you arent clear of?

OpenStudy (anonymous):

Isnt 189-49/4 176

OpenStudy (anonymous):

So where did you get 140 from?

OpenStudy (lxelle):

Remember your BODMAS.

OpenStudy (lxelle):

In this case i put bracket for (189-49) THERES A BRACKET. MEANS YOU GOTTA WORK OUT THE BRACKET FIRST.

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

So now when you do the brackets first why do you have to divide by 4 isn't that gonna give you a different answer?

OpenStudy (lxelle):

Noooo. I did the braacketed one first, I didnt divide the 4 to get 140. Only (189-49)= 140. Dont get confused.

OpenStudy (lxelle):

thats why you see i did \[\frac{ (189-49) }{ 40 }\]

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

So in the end the answer should be 35?

OpenStudy (lxelle):

yes

OpenStudy (lxelle):

My workings is just a guide for you. Try to work out your own workings. Youll understand better.

OpenStudy (anonymous):

Okay thanks I got it now. :)

OpenStudy (lxelle):

Okay. Glad that you got it! :)